OpenStudy (anonymous):

Please Help! Medal rewarded an isosceles trapezoid has an area of 40m^2 and an altitude of 2m. its two bases have an ratio of 2 is to 3. What are the lengths of the bases and one diagonal of the trapezoid.

OpenStudy (campbell_st):

well if you know the altitude is 2m then the area is 40 = h/2(a + b) where h = altitude and a and b are the parallel sides so 40 = 2/2(a + b) so really 40 = a + b the ratio is 2:3 so a = 2/5 of 40 b = 3/5 of 40 I'll leave you to do the calculations. hope it helps